OpenStudy (aonz):

Help please :) Find the values of a and b if the parabola y=a(x+b)^2 −8 has tangent y = 2x at the point P(4, 8),

OpenStudy (aonz):

i have gotten gradient to be 2a(x+b)

myininaya (myininaya):

ok great so far y' at (x,y)=(4,8) is 2a(4+b)=8a+2b And we are also given at that point we have y'=2x But that point is (4,8) so at that point we have y'=8 So this means we have 8=8a+2b Did you get this far?

myininaya (myininaya):

We should be looking for another liner equation

myininaya (myininaya):

Any ideas?

OpenStudy (aonz):

not really

OpenStudy (aonz):

i got 8=8a +2ab

myininaya (myininaya):

Well you also know (4,8) is on the parabola. See if you can use that.

myininaya (myininaya):

this should give you two equations with the only unknowns a and b

myininaya (myininaya):

You know how to solve a system of equations, correct?

OpenStudy (aonz):

err yea

myininaya (myininaya):

Did you think to do 8=a(4+b)^2-8

OpenStudy (aonz):

i got 8=8a +2ab not 8=8a +2b btw

myininaya (myininaya):

wait how did you get the first equation?

myininaya (myininaya):

oh wait i messed up when i wrote it

myininaya (myininaya):

2a(4+b)=8a+2b I didn't distribute there correctly 8a+2ab would be correct

myininaya (myininaya):

so yeah we have 8a+2ab=8

OpenStudy (aonz):

2nd equation is 16a + 8ba + ab^2 - 16 right?

myininaya (myininaya):

all of that = 0 yes

OpenStudy (aonz):

then we just solve right?

myininaya (myininaya):

yes

OpenStudy (aonz):

16a + 8ba + ab^2 - 16 = 0 and 8a+2ab-8= 0 err how?

myininaya (myininaya):

Did you try solving the linear equation either for a or b and replace the a or b in the first equation you ahve there

OpenStudy (aonz):

im pretty sure we made a mistake somewhere

OpenStudy (aonz):

the answers in the back of my book says 1/16 = a and 12= b

myininaya (myininaya):

why did you change the sign in that one equation?

OpenStudy (aonz):

oh typed it wrong before

OpenStudy (aonz):

oh wait...

OpenStudy (aonz):

can we start from the begining again?

OpenStudy (aonz):

the derivative is 8a+2ab that means the gradient is 8a + 2ab How did it become 8= 8a + 2ab

myininaya (myininaya):

y'=2x at (4,8)

myininaya (myininaya):

we also know that y'=2a(x+b) at (4,8) so at (4,8) we know 2x=2a(x+b)

myininaya (myininaya):

x=4 then

myininaya (myininaya):

so we have 2*4=2a(4+b)

myininaya (myininaya):

8=8a+2ab

OpenStudy (aonz):

ok alright

myininaya (myininaya):

your book's answer doesn't make sense

OpenStudy (aonz):

hmm

myininaya (myininaya):

so if a=1/16, b=12 and we have that y'=2x at (4,8) and y'=2a(x+b) at (4,8) then we have 2*4=2*1/16(4+12) correct? is this true?

myininaya (myininaya):

2*4=8 2/16*(4+12)=1/8*(16)=2 2 does not equal 8

OpenStudy (aonz):

err yea

myininaya (myininaya):

The equation you put is correct right? y=a(x+b)^2-8

myininaya (myininaya):

the point is (4,8) where we have the tangent line y=2x

OpenStudy (aonz):

yes right question

myininaya (myininaya):

i see nothing wrong with our resulting equations

OpenStudy (aonz):

hmm yea

OpenStudy (aonz):

but i dont think we could sub (4,8) into 16a + 8ba + ab^2 - 16=0

myininaya (myininaya):

we already did to get that equation

myininaya (myininaya):

remember y=a(x+b)^2-8 (4,8) is on that parabola

myininaya (myininaya):

8=a(4+b)^2-8

myininaya (myininaya):

a(4+b)^2-16=0 a(16+8b+b^2)-16=0 ab^2+8ab+16a-16=0

myininaya (myininaya):

So we have ab^2+8ab+16a-16=0 4=4a+ab

myininaya (myininaya):

Solve that bottom equation for either a or b then plug into the first equation.

OpenStudy (aonz):

a = 1, b=0

myininaya (myininaya):

By the way I just divided our earlier linear equation by 2 to get 4=4a+ab 4=a(4+b) 4/(4+b)=a Assuming b does not equal -4. Then we have 0=4/(4+b)*b^2+8*(4/(4+b))*b+16*(4/(4+b))-16

myininaya (myininaya):

Multiply both sides by 4+b 0=4b^2+32b+64-16(4+b)

myininaya (myininaya):

yep so i did make a mistake there the slope of 2x is 2 I put 8. haha So we had 2=8a+2ab was suppose to be that one linear equation not 8=8a+2ab